OpenStudy (anonymous):

show that determinant is equal to(b-c)(c-a)(a-b)(a+b+c),where R1=[1 1 1],R2=[a b c],R3=[bc ca ab]

OpenStudy (anonymous):

Now this you should really be able to do, as you only have to use Sarrus' rule. Look it up if you don't remember.

OpenStudy (anonymous):

The factorisation might be difficult though. Determinant = det [b c] - det [a c] + det [a b] [ca ab] [bc ab] [bc ca] = ab^2 - ac^2 - a^2b + bc^2 + a^2c - b^2c Considering (b-c)(c-a)(a-b) = (bc-ab-c^2+ac)(a-b) = abc - b^2c - a^2b + ab^2 - ac^2 + bc^2 + a^2c - abc = ab^2 - ac^2 - a^2b + bc^2 + a^2c - b^2c So answer is (b-c)(c-a)(a-b).

OpenStudy (anonymous):

plz solve it in easy way! i can not understand it.

OpenStudy (anonymous):

So I make an effort to solve your problem, and then you tell me to solve it again in the 'easy' way?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

plz solve my problem in sooo easy way because i solve that type questions at first time!